OpenStudy (anonymous):

solve by completing the square. p^2+3p-9=0 no answer yet tomas.a :)

OpenStudy (amistre64):

how do we complete a square?

OpenStudy (anonymous):

that's a good question

OpenStudy (anonymous):

you already did that in previous question didn't you?

OpenStudy (anonymous):

yes but this is tricky, it's my last one and i'm beat my friends

OpenStudy (anonymous):

(x+b)^2 = x^2+2b*x+b^2 you want the number b^2 to be on the end so you can factor it, you know what 2b is from the problem

OpenStudy (anonymous):

add (b/2)^2 to both sides

OpenStudy (anonymous):

-3-3sqrt5 over 2

OpenStudy (anonymous):

p^2+3p+2.25 = 9

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

p^2+3p+2.25 = 9 + 2.25

OpenStudy (anonymous):

(p+1.5)^2 = 11.25 p=+-Sqrt[11.25] - 1.5

OpenStudy (anonymous):

i'm sorry, none of what you are saying is making any sense.

OpenStudy (anonymous):

how do you foil (x+b)^2?

OpenStudy (anonymous):

tomas.a i got -3-3sqrt5over2

OpenStudy (anonymous):

yeah but don't foger plus/minus

OpenStudy (anonymous):

you are right, thanks!